# Hanse-Wissenschaftskolleg, Lehmkuhlenbusch 4, 27733 Delmenhorst, Germany

### v1

@@ 3.1 Linearization
Our aim now is to analyse the spectrum of linearized perturbations of these Q-balls. By analogy with the analysis of stability of the solutions in Smolyakov1 , we consider a perturbative expansion of the scalar field
@@
$\ddot{\xi}-\xi^{\prime\prime}+\left[V^{\prime}(f^{2})+V^{\prime\prime}(f^{2})f^{2}\right]\xi+V^{\prime\prime}(f^{2})\Phi^{2}\xi^{*}=0\,.$ (16)
@@ 3.2 Zero modes
However, apart from these two modes, there is another solution of the linearized equation (16) which satisfies the condition $\ddot{\xi}=0$: a stationary mode $\eta^{(Q)}(x)$ which transforms the frequency of the Q-ball $\omega$ as $\omega+\delta\omega$. Recall that the profile function $f(x;\omega)$ (10) depends on the frequency, therefore excitation of this mode affects both the frequency and the profile of the Q-ball, thus
$\eta^{(Q)}(x)=\partial_{\omega}f(x;\omega)+itf(x;\omega);\qquad(f(x;\omega)+\delta\omega\eta^{(Q)}(x))e^{i\omega t}\approx{f}(x;\omega+\delta\omega)e^{i(\omega+\delta\omega)t}\,.$ (25)
Note that both translational and phase shifting zero modes preserve the charge of the Q-ball while the frequency transforming mode affects it.
@@ Acknowledgements
PED was supported in part by the STFC under consolidated grant ST/T000708/1 “Particles, Fields and Spacetime”, and would like to thank the African Institute for Mathematical Sciences, South Africa, for hospitality during the last stages of this work. TR was supported by the Polish National Science Center, grant NCN 2019/35/B/ST2/00059. TR and DC were supported by the Priority Research Area under the program Excellence Initiative—Research University at the Jagiellonian University in Kraków. YS would like to thank the Hanse-Wissenschaftskolleg Delmenhorst for support.

### v2

@@ 3.1 Linearization
Our aim now is to analyse the spectrum of linearized perturbations of these Q-balls. By analogy with the analysis of stability of the solutions in Smolyakov1 ; Smolyakov2 , we consider a perturbative expansion of the scalar field
@@
$\ddot{\xi}-\xi^{\prime\prime}+\left[V^{\prime}(f^{2})+V^{\prime\prime}(f^{2})f^{2}\right]\xi+V^{\prime\prime}(f^{2})f^{2}\xi^{*}=0\,.$ (16)
@@ 3.2 Zero modes
However, apart from these two modes, there are two more solutions of the linearized equation (16), of the form $\xi(x,t)=\eta(x,t)e^{i\omega t}$ with the $\eta$ factor also time dependent: a stationary mode $\eta^{(Q)}(x,t)$, which transforms the frequency of the Q-ball $\omega$ as $\omega+\delta\omega$, and a mode corresponding to Lorentz symmetry. Recall that the profile function $f(x;\omega)$ (10) depends on the frequency, therefore excitation of this mode affects both the frequency and the profile of the Q-ball, thus
$\eta^{(Q)}(x,t)=\partial_{\omega}f(x;\omega)+itf(x;\omega);\qquad(f(x;\omega)+\delta\omega\eta^{(Q)}(x,t))e^{i\omega t}\approx{f}(x;\omega+\delta\omega)e^{i(\omega+\delta\omega)t}\,.$ (25)
The Lorentz symmetry mode is
$\eta^{(Lor)}(x,t)=t\partial_{x}f(x)+i\omega xf(x)\,,$ (26)
which follows from the expansion of a boosted Q-ball solution expanded up to $\mathcal{O}(v)$ terms:
$f(\gamma(x-vt))e^{i\gamma\omega(t-vx)}\approx\left[f(x)-v\left(t\partial_{x}f(x)+i\omega xf(x)\right)\right]e^{i\omega t}\,.$ (27)
Note that translational, phase shifting and Lorentz zero modes preserve the charge of the Q-ball while the frequency transforming mode changes it.
@@ Acknowledgements
PED was supported in part by the STFC under consolidated grant ST/T000708/1 “Particles, Fields and Spacetime”, and would like to thank the African Institute for Mathematical Sciences, South Africa, for hospitality during the last stages of this work. TR was supported by the Polish National Science Center, grant NCN 2019/35/B/ST2/00059. TR and DC were supported by the Priority Research Area under the program Excellence Initiative—Research University at the Jagiellonian University in Kraków. YS would like to thank the Hanse-Wissenschaftskolleg Delmenhorst for support. We would like to thank Mikhail Smolyakov for pointing out the presence of the Lorentz mode, which was missed in the previous version of the manuscript.